### v3

# IEEEexample.bib

Abstract
Methane ($CH_{4}$) is the second most prevalent greenhouse gas after carbon dioxide ($CO_{2}$), with a short atmospheric lifetime of approximately 9.14 years. Reducing methane emissions could quickly mitigate global warming. A methane plume is a localized concentration of methane. Methane emission monitoring involves three tasks: concentration inversion, plume segmentation, and emission rate estimation. The matched filter algorithm (e.g., Mag1c) for methane concentration inversion is sensitive to noise and reference signals. Research on methane plume segmentation is limited, often relying on subjective manual segmentation. Emission rate estimation typically uses the integrated mass enhancement (IME) algorithm, which is dependent on accurate meteorological measurements, limiting its applicability. Using PRISMA hyperspectral images from the WENT landfill in Hong Kong and the EnMAP hyperspectral dataset as a base map, a deep learning framework is proposed for quantitative methane emission monitoring from hyperspectral images based on physical simulation. We simulate methane plumes using large eddy simulation (LES) and create various fugitive methane concentration maps using the radiative transfer equation (RTE). We further augment the data to create a simulated EnMAP dataset $(https://github.com/Sigx093105/CH4_{p}lume_{d}ataset.git)$. We train a U-Net for concentration inversion, a Mask R-CNN for plume segmentation, and a ResNet-50 for emission rate estimation. These deep networks outperform conventional algorithms in validation accuracy and generalize well to RPISAM data. We combine tasks into multi-task learning models: MTL-01 and MTL-02, which outperform single-task models. This research demonstrates multi-task deep learning for quantitative methane monitoring and can be extended to various methane monitoring tasks.
Index Terms:
Methane plume, Hyperspectral remote sensing, Simulation, Multi-task deep learning.
I Introduction
Currently, landfills are still widely used around the world [magazzino2020relationship][zhang2019greenhouse]. It is estimated that about 40-50% of the greenhouse gases emitted by landfills are methane [bogner2007waste]. Landfills account for approximately 5% of global methane emissions and are considered one of the largest anthropogenic sources of $CH_{4}$ [stocker2013technical]. Furthermore, their contribution to atmospheric methane has been increasing since the beginning of the 21st century [canadell2021global]. The emission of greenhouse gases, including methane, not only affects the surrounding ecological environment [iravanian2020types] but also has the potential to exacerbate extreme weather events in urban areas [yang2021greenhouse]. From 1960 to 2019, the contribution of methane radiation stress accounted for 11% of total radiation stress, making it the second largest greenhouse gas after carbon dioxide. Compared to other main greenhouse gases, such as carbon dioxide and nitrogen oxides, methane exhibits a relatively short atmospheric life of approximately 9.14 years [canadell2021global]. This implies that reducing methane emissions may have a quicker impact on alleviating global warming compared to reducing emissions of other greenhouse gases.
To detect emitted methane plumes, airborne data (AVIRIS-NG) is commonly utilized due to small-scale methane emission sources [foote2020fast][foote2021impact]. However, the advent of PRISMA, a new-generation hyperspectral satellite with a spatial resolution of 30 m, has made it feasible to retrieve spaceborne methane emissions [cogliati2021prisma]. The monitoring process for methane plumes can be divided into three tasks: (1) methane concentration inversion, (2) methane plume segmentation, and (3) estimation of single plume flux rates. For methane concentration inversion, current methods predominantly rely on differential optical absorption spectroscopy (DOAS) and matched filter techniques. DOAS, although computationally intensive, is typically used for small-scale inversion in scenarios where emission point sources are known [thorpe2017airborne][thorpe2013high]. On the other hand, matched filter algorithms are commonly applied for large-scale methane plume screening [foote2021impact][thompson2015atmospheric]. However, it is important to note that the accuracy of the matched filter algorithm is heavily based on the precise modeling of the target spectrum and background estimation [foote2021impact][theiler2006effect01].
Few studies have been conducted on methane plume segmentation, but it is crucial to accurately delineate single plumes, as it significantly determines the accuracy of the estimation of the methane flux rate. Over the past few decades, a plethora of algorithms have been proposed for estimating point-source emission rates. However, most of these algorithms require local wind speed data as auxiliary information, which is difficult to obtain due to the limited availability of meteorological observation stations[varon2018quantifying] [jongaramrungruang2022methanet]. addressed this limitation by employing simulated plumes and deep learning techniques to estimate emission rates without the need for wind speed information, although the presented network could only estimate the emission rate of single-source plumes. In practical scenarios, it is often necessary to segment individual plumes from an image and estimate the emission rates separately.
Since the introduction of AlexNet in 2012, deep learning has witnessed rapid development and yielded remarkable results in various fields of remote sensing application[JIA202314][JIANG202330][HE202476]. The U-net deep learning architecture was initially designed for semantic segmentation tasks[ronneberger2015u][POLEWSKI2021297], but also exhibits enormous potential for regression[10.1145/3429309.3429328]. We regard methane concentration inversion as pixel-by-pixel regression, so U-net can be used to complete the task. Additionally, Mask R-CNN has proven to be highly effective in a wide range of instance segmentation tasks [SANIMOHAMMED2022100024] and should also be applicable to Methane plume segmentation. Moreover, ResNet addresses the issue of gradient vanishing, enabling deeper networks to improve feature extraction capabilities[he2016deep][shafiq2022deep]. Several earlier studies on emission rate estimation using shallow convolutional neural networks (CNN) have demonstrated feasibility, albeit with limited accuracy[9027926][guanter2021mapping]. It is reasonable to apply ResNet to improve the emission rate estimation. And our goal is to leverage physically simulated plumes as training data for multi-task deep learning techniques to simultaneously perform methane concentration inversion, methane plume segmentation, and flux rate estimation in a unified approach.
In this article, we start by describing the foundational principles of the algorithms utilized in our study, including large-eddy simulation, radiation simulation, the methane plume concentration inversion algorithm, the methane plume segmentation algorithm (instance segmentation), and the plume emission rate estimation algorithm. Following this, we provide a detailed account of our experimental setup, which encompasses the research area, the collected data, and the procedures used to build our methane monitoring dataset. Lastly, we evaluate the performance of deep learning algorithms against traditional ones for the tasks of methane plume concentration inversion, methane plume segmentation (instance segmentation) and plume emission rate estimation. Furthermore, we present the two multi-task learning networks, MTL-01 and MTL-02, that we developed for these tasks. The major contributions of this paper are as follows:
1. A physics-informed deep learning scheme is proposed for spaceborne retrieval of methane emissions based on data simulation using Large Eddy Simulation (LES), Radiative Transfer Equation (RTE), and data augmentation techniques.
2. An instance segmentation algorithm is applied to solve the problem of isolating and identifying methane emission sources by plume segmentation.
3. We show that the serialization of multiple sub-tasks for methane emission monitoring leads to additional error accumulation, and designed a multi-task learning model to suppress the errors.
II Method
In this section, we present the principles of methane plume simulation, methane signal simulation, and methane concentration inversion. The two multi-task deep learning networks and relevant training mechanisms that we have designed are also presented.
II-A Simulation
The simulation operation consists of two parts: simulation of methane plume emission and simulation of EnMAP-like satellite images with methane plume signal. For methane plume simulation, we utilized Large Eddy Simulation (LES), which is a computational fluid dynamics (CFD) method used to simulate turbulent phenomena in fluid flow[sagaut2002large]. It combines the advantages of Direct Numerical Simulation (DNS) and Reynolds-averaged Navier-Stokes equations (RANS) methods, aiming to accurately predict the statistical characteristics of turbulent flow[pope2001turbulent][domaradzki1997subgrid]. The core of LES lies in filtering the flow field into large- and small-scale vortices through a filtering operation. This can be achieved by applying a low-pass filter to remove high-frequency small-scale vortices while retaining the low-frequency large-scale vortices[domaradzki1997subgrid]. The filtered flow field equations form the basis of LES. The governing equations for LES are based on Reynolds-averaged Navier-Stokes equations (RANS)[germano1991dynamic], which are solved to simulate motion and turbulent effects in the fluid [10.1115/1.3124648]:
$\frac{\partial\bar{u}_{i}}{\partial t}+\frac{\partial}{\partial x_{j}}(\bar{u}_{i}\bar{u}_{j})=-\frac{1}{\rho}\frac{\partial\bar{p}}{\partial x_{i}}+\nu\frac{\partial^{2}\bar{u}_{i}}{\partial x_{j}\partial x_{j}}-\frac{\partial}{\partial x_{j}}(\bar{u}_{i}^{\prime}u_{j}^{\prime})$ (1)
where $\bar{u}_{i}$ is the time-average speed, $t$ is the temporal change in fluid motion, $x_{i}$ is the position of the fluid in a certain direction, $\bar{p}$ is the average pressure of a fluid at a location, $\rho$ is fluid density, $v$ is kinematic viscosity[chagovets2007effective], and $\bar{u}_{i}^{\prime}u_{j}^{\prime}$ is the reynolds stress[launder1975progress].
For simulation of EnMAP remote sensing images with methane plumes, we merged synthetic plume signals with real EnMAP remote sensing data[canadell2021global]. First, plumes of different shapes, sizes, and concentrations were generated by simulation and assigned corresponding spectral characteristics based on physical radiation properties. Then we fused the synthetic plumes with real EnMAP remote sensing images to generate simulated images with realistic background textures and synthetic methane spectrum. Such simulated images are crucial for training and evaluating our algorithms, as they can simulate real observation scenarios under controlled experimental conditions[foote2020fast]:
$L(\lambda){{}^{\prime}}=L(\lambda)*T_{plume}$ (2)
where $L(\lambda)$ denotes the real EnMAP image background spectrum and methane signal, $L(\lambda){{}^{\prime}}$ is the simulated spectrum, $T_{plume}$ is the simulated methane absorption cross-section. And $T_{plume}$ is generated by the summation of the radiative transfer equation (RTE), multiplied by the layered methane concentration and corresponding dry air density.
By combining these two steps, we can simulate methane plumes of various shapes, sizes, and concentrations and inject them into EnMAP data to generate simulated hyperspectral satellite images, which contain methane plume signals overlaid with a realistic scene background. This provides a solid foundation for our research and allows us to validate the performance and reliability of our methodology in real-world scenarios.
II-B Methane Concentration Inversion
Methane concentration inversion typically refers to a technique or process utilized to retrodict (or estimate) the concentration of methane on the earth’s surface or in the atmosphere from satellite images or other remote sensing data. This process involves the application of mathematical models and algorithms to transform satellite imagery into quantitative assessments of methane concentrations.
Mag1c is a matched filter method that effectively aligns trace gas concentration path lengths by incorporating sparsity and reflectivity correction. Mag1c is also one of the most advanced matched filter methods currently available and provides a mature and operational software interface[foote2021impact]. The calculation formula for Mag1c is as follows[foote2020fast]:
$a=\frac{(L_{i}-u)^{T}C^{-1}(t(u))}{(t(u))^{T}C^{-1}(t(u))}$ (3)
where $a$ is methane concentration, $L$ is radiation intensity, $t(u)$ is mean radiation, and $C$ is covariance matrix.
Methane concentration inversion in the West New Territories (WENT) region was performed using PRISMA data and the mag1c method. The results of this inversion were evaluated against a deep learning model to validate its generalization capabilities. Considering the intricate distribution of land and sea areas in Hong Kong, as well as the challenging rugged terrain and dense urban construction, it can be expected that the surface conditions in the region significantly influence the results of methane concentration inversion. To address this, we use a masking method based on K-means to run Mag1c[funk2001clustering].
In the realm of deep learning technologies, U-net is a prevalent and easily implementable architecture, frequently used for image semantic segmentation tasks[ronneberger2015u][POLEWSKI2021297][zang2021land]. Our rationale for selecting U-net as the model for methane concentration inversion comes from the following points: 1. We aim to achieve inversion of the methane concentration using a network that is as simplistic as possible. 2. Within the U-net framework, semantic segmentation is realized in a regression form, making it conveniently adaptable to the task of methane concentration inversion. 3. U-net can easily produce output images with the same length and width as the input image. Typically, the input and output channels of the U-net are of equal size. However, to accommodate the requirements for the inversion of methane concentration, an additional layer of $1*1$ convolution is added after the U-net encoding-decoding structure[lin2013network], and the loss function is replaced for the regression task. The loss functions used include the mean squared error (MSE) and $SmoothL_{1}$:
$loss_{mse}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y_{i}})^{2}$ (4)
$loss_{SmoothL1}=\begin{cases}\frac{1}{2}(y-\hat{y})^{2}&\text{if }|y-\hat{y}|<1\\ |y-\hat{y}|-\frac{1}{2}&\text{otherwise}\end{cases}$ (5)
where $y$ is the true label, $\hat{y}$ is the predicted value, $n$ is the number of samples, and $|y-\hat{y}|$ represents the absolute difference between $y$ and $\hat{y}$.
Before starting the training phase, various preprocessing techniques were applied to both labels and images. In particular, we computed the mean and standard deviation of all non-zero pixel values in the labels to achieve normalization. Referring to the pre-processing methods of the DOAS algorithm [wagner2004max][platt2008differential], the subsequent formula was utilized to preprocess the images:
${I_{after}=\frac{\log(I_{before})}{10}}$ (6)
where $I_{before}$ denotes the image prior to preprocessing, and $I_{after}$ denotes the same image post preprocessing. Both EnMAP and PRISMA have pixel values within the range of 0-10 after taking the logarithm, and dividing these values by 10 normalizes all preprocessed pixel values to fall within the range of 0-1. This technique helps to generalize the model across various sensors. The network is optimized using two loss functions, MSE and SmoothL1, starting with an initial learning rate of 0.001. The learning rate is decreased by 75% every 10 epochs, over a total of 200 epochs.
II-C Methane Plume Segmentation
Instance segmentation is a classical task in computer vision, and methane plume segmentation is essentially an instance segmentation task. Previous research work on methane plume segmentation is limited, and many studies are based on subjective segmentation criteria[gorrono2023understanding]. Some studies suggested using traditional algorithms for the segmentation of methane plumes, with contour tracking algorithms considered a good choice due to their efficiency and stability. On the other hand, there are many deep learning-based instance segmentation methods, among which Mask R-CNN is the most classical and widespread one[he2017mask]. It is a two-step instance segmentation algorithm that has proven to achieve high performance in a large number of datasets for a wide range of applications. Although new instance segmentation algorithms are continuing to emerge, some of the algorithms sacrifice segmentation accuracy for real-time operation[terven2023comprehensive], which is not necessarily required for our case. Other algorithms may achieve higher segmentation accuracy[wang2020solov2], but require an increasingly complex optimization process during training, which is not cost-effective. Therefore, Mask R-CNN is chosen as the deep learning algorithm for methane plume segmentation and active contour algorithms as the traditional counterpart. Typically, Mask R-CNN takes three image channels as inputs, but we modified the input layer to allow flexible alteration between 1 or 41 input channels to accommodate different segmentation requirements[foote2020fast], since the methane distribution map has only one band and EnMAP contains 41 bands after excluding the water vapor and carbon dioxide sensitive bands. The original loss function of Mask R-CNN is used:
$L=L_{cls}+L_{box}+L_{mask}$ (7)
where $L$ is the total loss function, $L_{cls}$ is the category loss function, which is based on the cross-entropy loss, $L_{box}$ is the bounding box loss function(As Mask R-CNN is a two-step instance segmentation model,its loss function inevitably includes the loss of object detection task), which is based on the $SmoothL_{1}$ loss, and $L_{mask}$ is the mask loss function, which typically uses the binary cross entropy:
$L_{cls}=-\sum_{i=1}^{N}y_{i}\log(\hat{y}_{i})$ (8)
where $y_{i}$ denotes the true label, $\hat{y}_{i}$ is the probability predicted by the model, and $N$ is the number of samples.
$L_{box}=\sum_{i=1}^{n}\text{Smooth}_{L1}(x_{i}-\hat{x}_{i})$ (9)
where $n$ is the number of bounding boxes considered in the current batch, $x_{i}$ represents a parameter set (including the center coordinates, width and height of the box) of the $i$-th predicted bounding box, $\hat{x}_{i}$ is the corresponding ground truth parameter for the $i$-th bounding box. During the training process, we use the minimum enclosing rectangle of the ground truth masks as the bounding box (for the predicted and ground truth bounding box).
$\text{Smooth}_{L1}(x)=\begin{cases}0.5x^{2}&\text{if }|x|<1\\ |x|-0.5&\text{if }|x|\geq 1\end{cases}$ (10)
where $x$ is the difference between the predicted value and the true value.
$L_{mask}=-\text{IoU}(b,\hat{b})+\frac{\rho^{2}(b,\hat{b})}{c^{2}}$ (11)
where $b$ and $\hat{b}$ respectively denote the ground truth binary mask and the predicted binary mask, $\text{IoU}(b,\hat{b})$ indicates the intersection over the union between these two binary masks, $\rho^{2}(b,\hat{b})$ is the Euclidean distance between the center points of two binary masks, and $c$ is a hyperparameter typically set to 10.
II-D Emission Rate Estimation
There are two scenarios for estimating emission rates. The first scenario involves estimating the emission rates using a supervised deep model with training samples collected specifically for this purpose. The goal here is to evaluate the performance of the estimation algorithm under ideal conditions. The second scenario refers to the estimation of emission rates based on plume segmentation results (MTL-01), where the objective is to evaluate and improve the accuracy of plume instance segmentation.
The emission rate of a methane plume is commonly estimated using the integrated mass enhancement (IME) mode[guanter2021mapping]. This mode involves the following calculations:
$Q=\frac{U_{eff}\cdot IME}{L}$ (12)
where $U_{eff}$ is the effective wind speed in $m/s$, $L$ is the plume length $m$, $IME$ refers to the integrated mass enhancement. And the calculations for $U_{eff}$, $L$, and $IME$ are as follows[varon2018quantifying][guanter2021mapping]:
$U_{eff}=0.34\cdot U_{10}+0.44$ (13)
where $U_{10}$ is the wind speed at 10 meters.
$L=\sqrt{A_{M}}$ (14)
where $\sqrt{A_{M}}$ is the area of a plume.
$IME=k\sum_{i=1}^{n_{p}}\hat{\alpha}(i)$ (15)
where $n_{p}$ is the number of pixels in a plume, $\hat{\alpha}(i)$ is the value of $ith$ pixel, and $k$ is a scale factor equal to $5.155\cdot 10^{-3}kg/ppb$.
ResNet is a deep learning method that uses residual blocks, which enables the neural network to learn the residual map by effectively avoiding issues such as gradient vanishing and explosion[he2016deep]. ResNet is widely applied in various application fields, including image recognition, object detection, and speech processing, yielding impressive results[veit2016residual][shafiq2022deep]. However, in this study, we pay particular attention to the application of ResNet in regression tasks, specifically for the complex and important task of emission rate estimation. Since the emission process is often affected by a variety of environmental factors and source dynamics, this makes emission rate estimation a highly nonlinear and complex problem. To meet this challenge, we chose ResNet-50 as our base network architecture for the following reasons: 1. Powerful feature extraction capabilities: ResNet-50 is able to learn and extract rich hierarchical features in an image through deep structure and residual connection, which is essential to capture subtle changes in methane emission from images. 2. Suitable for regression tasks: although ResNet-50 was originally designed for classification tasks, it can be easily adapted to regression tasks, such as emission rate estimation, by adjusting the output layer of the network (replacing the softmax layer with a linear layer).
We trained the ResNet-50 network with samples to perform the flux-rate estimation task. Subsequently, ResNet-50 was used as a pretrained network to validate the rationale of the plume segmentation results and assist in optimizing the instance segmentation network.
II-E Multi-task Learning I (MTL-01)
Experimental results indicated that the sequential prediction process involving Mask R-CNN followed by ResNet-50 led to substantial errors in emission rate estimation due to Mask R-CNN’s false positives and false negatives (Fig. 9). While these errors could be tolerable in the context of plume segmentation, we have developed a multi-task network named MTL-01 to tackle this issue. This network concurrently trains both Mask R-CNN and ResNet-50.
Considering that the loss function of Mask R-CNN measures instance segmentation and does not consider error propagation, we design a serial network (MTL-01), which takes the segmentation results of Mask R-CNN as input to ResNet-50 and calculates additional loss terms for back-propagation to assist in optimizing Mask R-CNN (Fig. 1). The ResNet-50 in the serial network is a pre-trained model, whose parameters are not updated. During the training process, the erroneous (over-)segmentation masks from Mask R-CNN are mainly categorized into two types. The first type is small false positive patches escaping from NMS(Non-Maximum Suppression) filtering[he2017mask], which were only found to slightly contribute to errors in subsequent emission rate estimation rather than main error source of the serial network. Simulation experiments indicate that the IME value of plumes is almost never below 300. Hence, we disregard false positive patches with IME values below 300 pixels, addressing this type of error using an object detection loss term. The second type of error consists of large false positive patches, often wrongly separated from overlapping plumes, causing substantial inaccuracies in emission rate estimation. To mitigate this, we introduce an additional loss term, $loss_{ER}$, for MTL-01. Although $loss_{ER}$ is still based on smoothL1, segmentation results with IME values below 300 pixels are excluded from its calculation. Missed masks are treated as having an emission rate estimate of 0 kg/h, and the loss is computed accordingly:
$loss_{ER}=\begin{cases}Smooth_{L1}(y_{i},\hat{y_{i}})&\text{if }y_{i}\in TP\\ Smooth_{L1}(0,\hat{y_{i}})&\text{if }y_{i}\in FP\end{cases}$ (16)
In MTL-01, we consider the $loss_{ER}$ term as a corrected loss term and assign it a coefficient lambda, while the original loss function of Mask R-CNN remains unchanged. Therefore, the loss function of MTL-01 is as follows:
$loss_{MTL\--{01}}=loss_{MaskR\,\--{CNN}}+\lambda\cdot loss_{ER}$ (17)
In this equation, $loss_{MaskR\,\--{CNN}}$ denotes the mask R-CNN loss function, $loss_{ER}$ denotes the loss function of emission rate estimation, and $\lambda$ is the coefficient. Based on our experiments, we found that $\lambda=0.1$ is optimal.
Figure 1: The structure of the multi-task learning model
The first multi-task learning model (MTL-01) is a sequential architecture combining a Mask R-CNN for plume segmentation and a ResNet-50 for estimating plume emission rates. The segmented plumes from Mask R-CNN are fed into ResNet-50 in real-time for emission rate loss calculation. The total loss for MTL-01 is the weighted sum of the individual losses of both networks. ’Multiplume’ describes a methane concentration map containing multiple plumes and noise. ResNet-50 is a 50-layer deep residual network. A Fully Convolutional Network (FCN) integrates semantic information from both low-level and high-level images via skip connections. In Mask R-CNN, ’heads’ are responsible for detecting objects and extracting masks. MLP stands for multi-layer perceptron. The second multi-task learning model (MTL-02) detects methane concentration and plume boundaries simultaneously. This model consists of a Mask R-CNN for plume segmentation and a U-Net for methane concentration inversion, both sharing a ResNet-50 feature extractor. While the Mask R-CNN’s decoding structure is unchanged, the U-Net decoding structure includes four skip connections, an ASPP module, and two 1x1 convolutional layers. ’Simulated EnMAP’ refers to EnMAP data augmented with simulated methane plume signals. The ASPP module comprises a 1x1 convolutional layer for channel reduction and multiple parallel atrous convolutional layers with adjustable dilation rates to capture features at different scales. ’Conv 1x1’ indicates a 1x1 convolution layer, ’Dilated Conv’ denotes an atrous convolution layer, and ’Pool’ refers to a pooling layer.
II-F Multi-task Learning II (MTL-02)
The process of methane concentration inversion entails estimating pixel-specific values using hyperspectral images, whereas methane plume segmentation is concerned with identifying plume masks within the same sets of images. Given their structural similarity and logical relationship, using multi-task learning models might improve the accuracy of both methane concentration inversion and plume segmentation tasks. Hence, a multi-task learning framework that merges U-net and Mask R-CNN architectures has been devised.
In MTL-02 (Fig.1), the mask R-CNN network architecture remains unchanged, and the U-net decoding structure is connected to the mask R-CNN backbone network. Considering the output sizes of the stages in the ResNet-50 backbone network, we added an additional output to the first convolution block of ResNet-50 to ensure that the size of the U-net decoding result matches the input size, similar to LinkNet[zhou2018d]. It should be noted that the deepest output of ResNet-50 is not involved in the construction of the U-net decoding structure but only participates in the FCN[long2015fully] structure of Mask R-CNN. To address the potential underfitting of the U-net decoding structure due to the max-pooling layer situated between two outputs of ResNet-50, we introduced an Atrous Spatial Pyramid Pooling (ASPP) layer directly after the U-net decoder. This ASPP layer consists of two dilated convolutions with dilation rates of 2 and 4. ASPP[chen2017deeplab], which builds on dilated convolutions and SPP (Spatial Pyramid Pooling)[he2015spatial], addresses the challenge of balancing the reduction in feature map resolution with the need to capture a large receptive field for image feature extraction. ASPP leverages dilated convolutions with varying dilation rates to encapsulate the image context at multiple scales. In this approach, the input feature map is subjected to several dilated convolutions with different rates, which are then combined. To further expand the model’s receptive field, the resulting feature maps from these dilated convolutions are subjected to average pooling.
In the loss function for $MTL-02$, a traditional approach is employed to use a weighted average of the individual loss functions for multiple tasks[bruggemann2021exploring]:
$loss_{MTL\--{02}}=w_{1}\cdot loss_{U\--{net}}+w_{2}\cdot loss_{MaskR\,\--{CNN}}$ (18)
where $loss_{U{-}net}$ represents the loss function of U-net, $loss_{Mask\,R{-}CNN}$ denotes the loss function of Mask R-CNN, and $w_{1}$ along with $w_{2}$ are the respective weights for these loss functions.
To determine the weights for the MTL-02 loss function, we utilized the Dynamic Weighted Average (DWA) algorithm[liu2019end]. The main idea is that tasks differ in their difficulty and speed of learning. Instead of giving each task a separate learning rate, it is better to adjust the weights dynamically so that all tasks advance at a similar pace. In simple terms, tasks that show a faster decrease in loss receive smaller weights, while those with a slower decrease are allocated larger weights.
${w}_{k}(t-1)=\frac{L_{k}(t-1)}{L_{k}(t-2)}$ (19)
${\lambda}_{k}(t)=\frac{K\cdot e^{\frac{{w}_{k}(t-1)}{T}}}{\sum_{i=1}^{K}e^{\frac{{w}_{i}(t-1)}{T}}}$ (20)
where $K$ is the number of tasks, $\lambda_{k}(t)$ is the weight of the task k at the i-th iteration, and $L_{k}(t)$ is calculated as the average loss in each epoch over several iterations. When $T$ is a temperature, $T=1$ indicates that $\lambda$ is equal to the result of softmax; However, as $T$ increases significantly, $\lambda$ approaches 1, leading to equal loss weights for all tasks.
III Experiments
III-A Research area and data
III-A1 Research area
Situated in the southern part of China, Hong Kong is predominantly marked by hilly terrain. The region, which is highly urbanized, experiences a humid and rainy climate, with an average annual precipitation exceeding 2000 millimeters. Within Hong Kong, there are two key landfills: West New Territories (WENT) landfill and North East New Territories (NENT) landfill. Since most of the PRISMA images were obtained from the WENT region, our generalization verification is mainly concentrated in this area. The Tuen Mun landfill (Fig. 2), which spans approximately 27 hectares, is located in the southwest of the Tuen Mun district near Shenzhen Bay. It is one of the limited waste disposal sites sanctioned by the government in the Hong Kong Special Administrative Region and primarily receives municipal solid waste from various areas, including Hong Kong Island, Kowloon and New Territories.
III-A2 Hyperspectral Data
The PRISMA satellite was launched on 22 March 2019, and offers hyperspectral images of global coverage with a spatial resolution of 30 m. The spectral smile is less than 5 $nm$, the spectral resolution is better than 12 $nm$ in a spectral range of 400-2500 nm (VNIR and SWIR regions) [cogliati2021prisma], and datasets are open-access. We have collected several PRISMA images of WENT, which are: $20220102$, $20220119$, $20220914$, $20230217$, $20230301$, $20230307$, $20230416$. These raw images will be used to verify the generalizability of deep learning models[thorpe2013high][foote2020fast].
Figure 2: Research area
The research site is the WENT landfill situated in the Tuen Mun District of Hong Kong SAR. The landfill is located on a hillside close to Shenzhen Bay, as indicated by the blue rectangle on the map.
[…]
III-B Generation of diversified plumes in terms of emission rates and wind speeds
The simulated plumes were generated using Palm software[2001PALM][maronga2015parallelized],using the average atmospheric pressure at sea level in Hong Kong (0.96 hPa) and a temperature of 296K. A specialized smog mode was implemented, including a methane dispersion feature, to simulate the diffusion of methane point sources within the troposphere from 0 to 10,000 meters. We defined four gradients for the methane point source emission rate: 500kg/h, 1000kg/h, 1500kg/h, and 2000kg/h. The wind direction was set from west to east, with wind speed gradients of 1m/s, 2m/s, 3m/s, 4m/s, 5m/s, 6m/s, 7m/s, 8m/s, 9m/s, and 10m/s (Fig. 3). For each wind speed gradient, the methane point source emission rate gradients were simulated in sequence. The simulation period covered 0 to 2.5 hours, with snapshots taken at 30-second intervals from 1 to 2.5 hours, resulting in a total of 7200 plume snapshots. Among them, the simulated plume emission data with wind speeds of $2m/s$, and $10m/s$ were used to create the validation dataset, the simulated plume emission data with wind speeds of $1m/s$, and $9m/s$ were used to create the test dataset, while the remaining plume emission data were used to create the training dataset.
Figure 3: Snapshot of a methane plume simulated by the Palm software
This figure illustrates the methane plume at various emission rates and wind speeds, modeled with the Palm software. Each image was taken 180 minutes after the initiation of the emission, with simulation parameters specified in Section II.B.
III-C Generation of multi-task learning datasets
The dataset we are going to create consists of three subsets: inversion dataset, segmentation dataset, and emission rate dataset. Each subset comprises three groups of data, namely Train, Val, and Test, which are respectively used for training, validation, and testing. Each group of data contains several samples, and each sample includes an image and its corresponding label. We assign plume snapshots of different wind speed gradients to Train, Val, and Test. The base map is also randomly divided into three groups and allocated to Train, Val, and Test accordingly. The creation process of a single sample group is as follows(Fig. 4):
By capturing an image of a plume and averaging the methane concentrations at various heights, a 2D representation of a single plume with an emission rate of $m$ kg/h is produced. Scaling all pixel values in this 2D plume by a factor of $a$ results in a new emission rate of $m\times a$ kg/h, which is used as the sample label in the emission rate dataset. To ensure plume diversity, we randomly generate a threshold between 0.05 and 0.10, and set pixels in the 2D plume below this threshold to zero. A rotation enhancement ranging from -170° to 170° is then applied to the 2D plume, creating the image counterpart for the emission rate dataset sample. Plume segmentation is treated as an instance segmentation task. Initially, we create a blank $256\times 256$ image filled with zeros. We then randomly select $N$ images from the emission rate dataset (where $N\leq 3$ and $N$ can be 0) and superimpose them onto the all-zero image to form a multi-plume image. Concurrently, we determine the boundaries of each plume and assign corresponding labels to the segmentation dataset. If $N=0$, the dataset will lack labels. Lastly, we randomly generate $256\times 256$ Gaussian white noise and add it to the multi-plume image[jongaramrungruang2022methanet][gil2002efficient] to create the image for the segmentation dataset. Additionally, the overlap ratio between any two plumes must not exceed 15
$Overlapratio(A,B)=\frac{\sum_{x_{i}\in A\cap B}x_{i}}{\sum_{x_{i}\in A\cup B}x_{i}}$ (21)
where A,B represent two 2D plumes, $x_{i}$ is the pixel value in the plume.
The inversion data originate from multi-plume images created during the process of generating the segmentation dataset. These multi-plume images serve as labels for the inversion dataset samples. First, a multi-plume image is obtained and inspected for any emptiness. If found to be empty, a random EnMAP image is chosen as the sample image. On the other hand, if the multi-plume image is not empty, methane transmission spectra for each pixel value are constructed on a column-by-column basis. For columns that are entirely zeros, the transmission spectrum is defined as a $256\times 1\times 41$ matrix where all elements are set to 1. For columns with non-zero values, an optical depth matrix with dimensions $256\times 1\times K$ (where $K$ is the number of elements in the absorption cross-section dataset) is created using the approach explained in Section II.A. The negative logarithm of the pixel values in the optical depth matrix is then used to compute the original transmittance matrix.
First, the central wavelengths and half-bandwidths of the 41 EnMAP bands are determined to construct the spectral response function (SRF)[green1998spectral][2008Toward]. Then, the SRF is utilized for normalized convolution along the $K$ dimension of the initial transmittance matrix, resulting in a $256\times 1\times 41$ ’column’ transmission spectrum matrix. Following the processing of each column of the non-empty multi-plume image, all the transmission spectrum matrices are combined sequentially in their original column order to form a $256\times 256\times 41$ transmission spectrum matrix.
In the final step, the transmission spectrum matrix is merged with random EnMAP data through the Hadamard product to produce the image for the non-empty multi-plume. It’s important to note that, since the simulated methane plumes are mainly below 3000 meters in altitude, we use the average temperature and atmospheric pressure of Hong Kong to determine the methane absorption cross-section and calculate the optical depth based on the mean atmospheric density from 0 to 3000 meters. While this method may reduce some accuracy in spectral simulation, the compromise is worthwhile given the substantial computational load of creating the inversion dataset. The entire code for generating the inversion dataset is implemented in PyTorch and executed on an NVIDIA GeForce RTX 3060 GPU.
Figure 4: Generation process of training samples
The training dataset is composed of three types of samples: Emission rate sample, Segmentation sample, and Inversion sample. The training dataset consists of numerous examples of these types of samples. (a) A snapshot of a plume, simulated by Palm, is scaled (including the emission rate), and concentration values below a certain threshold are set to zero. A random rotation from -170° to 170° is applied to produce a single enhanced plume, which acts as the emission rate sample image. The scaled emission rate serves as its label. (b) Up to N ($N\leq 3$) plumes generated in step (a) are randomly merged (ensuring that overlap does not exceed 70%) to create a multi-plume image. Noise from mag1c inversion outcomes is added to this multi-plume image to generate the segmentation sample image, with the boundary of each plume noted as the label. (c) Following the method in (a), a transmission spectrum matrix that matches the EnMAP base map dimensions is generated based on the multi-plume image from (b). The simulated EnMAP image is then created by applying the Hadamard product of the transmission spectrum matrix and the EnMAP base map, which serves as the inversion sample. The multi-plume image serves as the label for the inversion sample.
III-D Accuracy Assessment
III-D1 Regression Task
Both inversion of methane concentration and estimation of emission rate are integral to the regression mission. Hence, to evaluate accuracy, we employed the RMSE and MAE metrics.
$RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y_{i}})^{2}}$ (22)
$MAE=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y_{i}}|$ (23)
where $y$ is the ground truth value, $\hat{y}$ is the predicted value, and $n$ is the number of samples.
III-D2 Segmentation Task
Average precision (AP) is used to measure the accuracy of the methane plume segmentation task. The formula for calculating AP is as follows[he2018hashing]:
1. Set a threshold and determine the number of $TP$, $FP$, $TN$, and $FN$ according to the threshold.
$TP$: True positive, the number of positive samples correctly classified as positive.
$FP$: False positive, the number of negative samples falsely classified as positive.
$TN$: True negative, the number of negative samples correctly classified as negative.
$FN$: False negative, the number of positive samples falsely classified as negative.
2. Calculate precision by dividing the number of true positives by the sum of the number of true positives and false positives.
$Precision=\frac{TP}{TP+FP}$ (24)
3. Calculate the recall by dividing the number of true positives by the sum of the number of true positives and false negatives.
$Recall=\frac{TP}{TP+FN}$ (25)
4. Repeat this process for different thresholds to obtain a set of precision and recall values.
5. Calculate the AP value:
$AP=\sum_{i=1}^{n-1}(R_{i+1}-R_{i})P_{inter}(R_{i}+1)$ (26)
where $(R_{i+1}-R_{i})$ represents the span of the recall interval, and $P_{inter}(R_{i}+1)$ denotes the interpolated precision at a recall of $R_{i}+1$. The AP value is computed by aggregating all interpolated precision values, each weighted by their respective recall interval lengths between consecutive recalls.
IV Results and Discussions
IV-A Simulation
To verify the accuracy of simulated plumes, we consulted relevant studies[gorrono2023understanding]. The integrated methane emission (IME) values for a typical methane plume should exhibit a linear relationship with its emission rate. We computed the IME values for all 2D plume snapshots, plotted them against respective emission rates in a scatter plot, and fitted a linear curve, from which the coefficient of determination is derived. The findings show that the coefficient of determination for the linear curve was 0.88 (Fig. 5), indicating a linear correlation between the IME values and emission rates of these 2D plume snapshots. Therefore, we conclude that the quality of simulated methane plumes is reasonable to a large extent.
Figure 5: Correlation between emission rates and simulated plume dataset IME values.
We analyzed all 2D plume images generated by the PALM software and determined their IME values. A scatter plot was created to display the IME values of the plume images against their corresponding emission rates, and a linear fit was applied. The findings indicate a strong correlation between the IME values and emission rates of these plume images.
We utilize simulation methods to create EnMAP datasets that include plume signals. Although the production process follows theoretical principles, the physical model is a simplified approximation of real-world conditions. Therefore, it is crucial to verify the accuracy of incorporating plume signals into the EnMAP base map. Without actual labels, we apply an end-to-end approach to validate the simulated EnMAP images. First, a random 2D plume snapshot is taken and a transmittance matrix is created following the method described in Section II. Next, a random EnMAP base map is chosen and combined with the transmittance matrix using the Hadamard product to generate a simulated EnMAP image. Following this, the Mag1c tool is employed to extract the methane concentration distribution map from the simulated EnMAP image. Finally, we compare the 2D plume snapshot with the methane concentration distribution map. If their methane concentration distributions match closely, the plume signal incorporation process can be deemed valid. In Fig. 6, the left image shows a 2D plume snapshot, while the right image presents the methane concentration distribution map derived by Mag1c from the relevant EnMAP data. It is notable that the right image contains some additional noise compared to the left image, but the primary features are quite similar, thereby validating the plume signal incorporation process.
Figure 6: The results of end-to-end verification for simulation.
Due to the absence of the ground truth value of methane concentration, the evaluation is performed through an end-to-end method: comparing the methane plume extracted from the simulated EnMAP image using mag1c with the methane plume added to the EnMAP base map. The results show that the methane plume extracted by mag1c closely resembles the plume incorporated into the EnMAP base map, indicating that the addition of the plume signal is appropriate.
IV-B Methane Concentration Inversion
Each model was assessed on the test dataset for methane concentration inversion (see Table I). The results from the validation suggest that the deep learning model consistently surpasses Mag1c in inversion accuracy, with U-net employing MSE yielding better precision than with SmoothL1. Reviewing the comprehensive validation results, it can be deduced that the U-net network exhibits superior predictive capabilities compared to Mag1c. However, the validation using simulated data may introduce some bias. To address this, we further evaluated U-net’s generalization performance using PRISMA data from WENT.
TABLE I: VALIDATION RESULTS OF METHANE CONCENTRATION INVERSION
Methods RMSE$/ppm$ MAE$/ppm$
Mag1c 0.5377 0.1392
U-net + Smooth $L_{1}$ loss 0.1208 0.0225
U-net + MSE loss 0.0926 0.0159
MTL-02 0.0790 0.0196
After eliminating the sensitive bands for water vapor and carbon dioxide, PRISMA and EnMAP retained 49 and 41 bands, respectively. The PRISMA data were then resampled to 41 bands in the spectral dimension, aligned with the center wavelengths of the 41 EnMAP bands used in the prediction. Subsequently, the pre-trained U-net was applied to invert the methane plume in the area, and its results were compared to the Mag1c inversion based on the 49 PRISMA bands. As depicted in Fig. 7, three representative inversion visualizations were presented. It is observable that U-net and Mag1c provide consistent inversion results in certain datasets, while some differences are evident in others. Due to the lack of true methane plume concentration values in the WENT region, determining the most accurate method is challenging. We investigated potential reasons for inconsistencies between the two approaches: 1. The uneven terrain of the WENT region and proximity to the ocean create a complex noise distribution in the PRISMA data, affecting the methods differently. 2. The PRISMA data resampling process results in some plume information loss, leading to disparities in U-net and Mag1c inversion results. 3. The limited number of EnMAP base maps used for U-net training in rugged areas may cause U-net to predict methane column concentrations based on smoother terrain.
Figure 7: Comparison of generalization results between U-net and Mag1c on real data.
Sample illustrations of U-net’s generalization performance on PRISMA image.
IV-C Methane Plume Segmentation
The Mask R-CNN model was trained on two versions of the plume segmentation dataset as detailed in Section II.C. One version used simulated hyperspectral images (hyperspectral image) as input, while the other used methane plume distribution maps (single). Despite this, the labels for both versions of the data set were identical. Similarly, model validation was performed separately using validation sets corresponding to the two versions of the data set, maintaining identical labels. Note that the Active Contour method was validated exclusively using single methane plume maps as input because it is an unsupervised image segmentation technique that requires an explicit relationship between pixel values and task-dependent physical properties. The results indicated that the Mask R-CNN model achieved higher precision than the Active Contour method (Table II). However, when assessing the AP metric at $AP_{95}$, the Active Contour method outperformed Mask R-CNN, suggesting that Active Contour excels in isolating individual plume masks. Despite this, the overall performance of Active Contour is inferior to that of Mask R-CNN due to its difficulty in effectively distinguishing overlapping plumes and its tendency to create small, extraneous segmented areas.
TABLE II: VALIDATION RESULTS OF METHANE PLUME SEGEMENTATION
Models $AP_{50}$ $AP_{75}$ $AP_{95}$ $AP_{50:95}$
Active Contour 42.03 39.04 9.35 34.96
MTL-01 84.06 79.93 0.00 64.74
Mask R-CNN single 82.58 76.73 0.00 61.84
Mask R-CNN hyper 29.28 14.70 0.00 15.07
MTL-02 35.21 15.44 0.00 17.14
Both versions of Mask R-CNN demonstrate superior precision in extracting the plume mask from the methane plume distribution map compared to extracting it directly from hyperspectral images. This is primarily due to two factors: first, the EnMAP data, produced through a two-step simulation process, inherently contain more errors than the one-step simulation method. Second, extraction of the EnMAP plume mask involves both inversion and segmentation steps, which increases the likelihood of inaccuracies compared to direct extraction from the methane plume distribution map. Regarding the ”single” Mask R-CNN, despite its capability to efficiently segment overlapping plumes, its segmentation outcome still exhibits some errors (Fig. 8). These errors can generally be classified as follows: 1. If a plume is divided into two sections, the single Mask R-CNN may incorrectly segment it into two separate plumes. 2. For overlapping plumes, the single Mask R-CNN may generate redundant or duplicate segmentations.
Figure 8: Segmentation results of various methods on the test dataset
The region outlined by the blue border shows the segmentation of methane plumes produced by different methods.
To evaluate the Mask R-CNN model’s ability to generalize to real-world contexts, mag1c is used to reverse-engineer methane concentration maps from various PRISMA images within the WENT area. Subsequently, both Active Contour and Mask R-CNN were applied to delineate methane plumes. The blue boxes in Fig. 9 denote these identified plumes. It is evident that Active Contour maps all connected pixel regions, whereas Mask R-CNN typically highlights funnel-shaped pixel regions, indicating some resistance to noise. Despite this, both methods identified a comparable number of methane plumes. We acknowledge the absence of precise data on the exact count of methane emission sources (or plumes) in the WNET area. Nevertheless, visual inspection reveals that both techniques detected the major plumes and overlooked several smaller ones on the right. This omission could be attributed to the rugged coastal terrain of the WENT region, where plume dispersion is influenced by intricate factors like topography and sea-land breeze dynamics, resulting in more complex plume formations.
Figure 9: Comparison of generalization results of different methods on real plume images
To assess the generalizability of the Mask R-CNN model on real-world data, we employed mag1c to invert the methane concentration map derived from multiple PRISMA datasets in the WENT region. Subsequently, we used both Active Contour and Mask R-CNN to extract the methane plume. The area marked by the blue line in the figure indicates the extracted methane plume. It is evident that Active Contour aims to extract all connected pixel regions extensively, while Mask R-CNN focuses on pixel regions with a funnel-shaped distribution. Nonetheless, the overall count of methane plumes identified by both methods is comparable.
IV-D Estimation of Methane Emission Rate
We employed ResNet-50 and AlexNet to estimate emission rates by training them on the provided dataset and evaluating the accuracy of their estimations on the corresponding test set (Table III). The results reveal that ResNet-50 outperforms the IME algorithm in the validation set. Although AlexNet’s estimation accuracy is slightly higher than that of the IME algorithm, it still does not match ResNet-50’s performance. This aligns with the common understanding in deep learning that deeper networks enable each layer to address simpler tasks. As a result, more complex nonlinear representations are possible, which helps to learn intricate mappings[2016On], and boosts the generalization ability and robustness of the network.
TABLE III: VALIDATION RESULTS OF EMISSION RATE ESTIMATION OF SINGLE NETWORKS
Methods RMSE$/kg*h^{-1}$ MAE$/kg*h^{-1}$
IME 187.47 142.05
AlexNet + MSE loss 177.31 141.03
ResNet + MSE loss 126.93 96.09
To assess ResNet-50’s ability to generalize to real-world data, we employed Mag1c to invert methane concentration distribution maps derived from RPISMA data in the WENT region and manually identified 28 methane plumes. We estimated their emission rates using both the IME method and the pre-trained ResNet-50 model, constructed scatter plots, and performed linear curve fitting. As shown in Fig. 10, the scatter plot illustrates that the majority points are aligned closely to a straight line, indicating that the predictions of both approaches are typically in good agreement with the real-world dataset. However, it is also apparent that the ResNet-50 model shows a tendency to saturate in areas with lower emission rates.
Figure 10: Comparison of generalization results between ResNet-50 and IME on real data
We manually extracted 28 methane plumes from the inversion results of PRISMA’s WENT region and predicted their emission rates using the IME method and the trained ResNet-50 model separately. The prediction results are shown in the figure. Both methods yielded comparable results. However, it can also be observed that ResNet-50 shows some saturation in regions with low emission rates.
IV-E Serial Network
Although individual models demonstrated high accuracy for each specific task, the efficacy of serializing multi-network models still warrants further examination. Initially, we combined Mask R-CNN and ResNet-50 to handle the tasks of segmenting methane plumes from concentration maps and estimating their emission rates. We assessed the accuracy of these emission rate estimates in the validation set under three scenarios: 1. For true positive segmented plumes, we used pre-trained ResNet-50 to estimate emission rates and compared them with the label’s emission rates. 2. For false-positive segmented plumes, we similarly used ResNet-50 to estimate their emission rates and compared them against 0 kg/h. 3. For misdetected plumes, we assumed their emission rate estimates to be 0 kg/h and compared these estimates with the corresponding labels. Ultimately, RMSE and MAE between all predicted emission rates and label values were calculated (Table IV). The results indicated that the accuracy of the emission rate estimation using Mask R-CNN + ResNet-50 was significantly lower than using ResNet-50 for single-task emission rate estimation. It indicates that the serialization of Mask R-CNN and ResNet-50 could introduce additional errors in the overall performance of methane flux rate estimation.
TABLE IV: VALIDATION RESULTS OF EMISSION RATE ESTIMATION OF SERIAL NETWORKS
Methods RMSE$/kg*h^{-1}$ MAE$/kg*h^{-1}$
ResNet-50 + MSE loss 126.93 96.09
Mask R-CNN+ResNet $plume_{300}$ 1157.45 859.56
MTL-01 1038.30 788.31
To investigate additional sources of error, we chose segmented plumes with significant errors (e.g., exceeding 500 kg/h). The analysis identified two additional error sources: 1. Errors stemming from small patches that are essentially noise; 2. Errors arising from false positive patches.
We examined the distribution of minor noise patches and discovered that the majority are defined by a mask size of less than 300 pixels, while authentic plume mask sizes seldom fall below 300 pixels. As a result, we exclude segmented masks smaller than 300 pixels in size and reassess the accuracy of Mask R-CNN + ResNet-50 for estimating emission rates.
IV-F Multi-task Learning I (MTL-01)
We trained the $MTL-01$ network, and after 30 epochs, $MTL-01$ outperformed Mask R-CNN in the plume segmentation task. As a result, the loss in estimating the emission rates of multiple methane sources was also lower than when using a serialized approach with $Mask\,R-CNN$ and $ResNet-50$. Additionally, by simulating a smaller training dataset with both methods, $MTL-01$ demonstrated even higher accuracy compared to the serialized training of $Mask\,R-CNN$ and $MTL-01$. It is evident that the separation of extra plumes is suppressed (Fig. 9), which lowers the predicted emission rate. $MTL-01$ mainly segments the core regions of the jet stream while overlooking some smaller sections of the tail. This has led to improvements in both the plume segmentation task and the emission rate estimation (Table II, Table IV).
Figure 11: Comparison of the sequential network and multi-task learning network I (MTL-01).
IV-G Multi-task Learning II (MTL-02)
We trained the MTL-02 model, which reached convergence after 30 epochs, showing a rapid decline in both loss function values and validation accuracy values. In the end, MTL-02 obtained slightly better accuracy in methane concentration inversion than U-net (Table I), and superior accuracy in plume segmentation compared to Hyper Mask R-CNN (Table II). This supports the notion that multi-task learning could enhance the convergence rate and fitting accuracy of the model, likely due to the shared information among various labels.
IV-H Limitations
Monitoring methane plumes is an intricate issue, and the current research in this domain is neither thoroughly explored nor systematic. Our proposed method also has some constraints: 1. The shapes of the simulated plumes that we used mostly adhere to a standard Gaussian model, which represents an ideal scenario. However, in reality, methane emissions from point sources are influenced by a variety of factors like topography and humidity, leading to non-Gaussian plumes. Whether our method applies to non-Gaussian plumes remains an open question; 2. The maximum overlap ratio of the plumes that we examined is $15\%$. However, naturally, many sources of methane emission are situated close to each other, and the overlap ratio of two plumes can be significantly higher than $15\%$. In such scenarios, even though we can generate highly overlapped methane plumes, it is challenging to provide precise training labels, necessitating the exploration of point-based semi-supervised learning techniques to achieve reasonable segmentation outcomes for these highly overlapped plumes; 3. The acquisition of hyperspectral satellite imagery is significantly affected by weather conditions, and hyperspectral image data sources are scarce in low-latitude coastal regions. Hence, a solution relying solely on hyperspectral images may not fulfill high-frequency monitoring needs. It is essential to devise an ensemble estimation method based on multiple data sources.
V Conclusion
Currently, there are significant uncertainties in monitoring fugitive methane emissions using satellite imaging spectrometers. We have created a demonstrative method utilizing deep learning models and simulated hyperspectral images for methane concentration inversion, plume segmentation, and emission rate estimation in landfills. Specifically, U-net was employed for methane concentration inversion, yielding higher accuracy and reduced noise compared to mag1c. Mask R-CNN was used for plume segmentation, eliminating subjectivity in the process. For estimating methane plume emission rates, ResNet-50 is utilized to bypass the dependence on wind speed measurements. We also addressed additional errors in emission rate estimation caused by serial networks by developing a multi-task model (MTL-01) to reduce these errors and enhance the precision of plume segmentation. Furthermore, a multi-task model (MTL-02) that performs both methane concentration inversion and plume segmentation was proposed, achieving higher validation accuracy than individual single-task models. Finally, we highlighted the limitations of the current approach, emphasizing the need for future research on highly overlapped plumes, non-Gaussian plumes, and the integration of multi-source data for inversion.
Acknowledgment
The work described in this paper was substantially supported by a grant from the ECF and EPD of the Hong Kong Special Administrative Region, China (Project No. PolyU ECF111/2020).
[…]